# On the composable security of weak coin flipping

### v2

@@ 3.2 Impossibility for a general resource
The composable security is defined by comparing between the ideal resource and the actual protocol. Theorem 5 denies the possibility of constructing an actual protocol close to the ideal WCF resource in Definition 4. However, if we modify the definition of the ideal WCF resource into a presumably more complicated one, then there might still exist some protocol that constructs this more complicated resource.
@@
  • Simple. By simplicity of resources, we typically refer to the process of abstracting away specific details of a resource that reduces its overall interaction complexity and allows to study it in a more general context. Here we take the interaction complexity $\mathsf{Intcom}(S_{A})$, $\mathsf{Intcom}(S_{B})$ (defined in Section 2.2) to quantify it.
Next, we will show that a fully expressing and simple WCF resource cannot be constructed from noiseless communication channels. Specifically, Theorem 6 states that, if a fully expressing resource for $\mathrm{WCF}(z,\epsilon)$ with small $\epsilon$ can be constructed from noiseless communication channels, then its interaction complexity cannot be low.
@@
To prove Theorem 6, we need a proposition rewritten from [40, Theorem 8.2].
@@
Some parts of the argument in the above proof are of more general interest. That is, if there exists a resource $\mathcal{S}$ that abstracts a two-party protocol $\pi$, it implies the existence of another protocol $\pi^{\prime}$ with the same interaction complexity as the resource $\mathcal{S}$. Furthermore, if the resource reduces the interaction complexity of $\pi$, then $\pi^{\prime}$ is a better protocol than $\pi$ in terms of interaction complexity. An example is delegated quantum computation (DQC), where the client aims to delegate computation to the server while maintaining privacy. The $S_{A}$ (against dishonest client) of DQC resource [41] abstracting the protocol in [42, 43] has no interaction in either side, because it is just trivial to talk about a dishonest client. This implies the existence of a simple DQC protocol with no interaction, wherein the client performs all the computation herself.
@@ 4 Security of weak coin flipping under composition
Section 3 highlights the difficulties when formulating the security of WCF in the AC framework. Fortunately, despite these difficulties, there is a prospect of attaining positive outcomes regarding the composability of such protocols. This section starts with the limitation of the stand-alone security definition (Definition 3) in the context of composition. Specifically, for the construction of unbalanced WCF from a balanced one, we propose a condition (Inequality (26)) which, while implicit, is essential in the security proof of the unbalanced WCF protocol depicted in [29]. Subsequently, we introduce global security as an extension of the stand-alone definition to encompass possible interdependencies between the outputs of the WCF protocol and other components in a larger system. Furthermore, we establish the condition under which global security is maintained.
@@ 4.1 Stand-alone security is not sufficient
For a cryptographic task, some security definitions seem fine from a stand-alone viewpoint but may present some loopholes when composed with other protocols. For example, the accessible information criteria for quantum key distribution falls short of providing composable security due to the data-locking effect[2, 1]. This issue is similar to WCF. In particular, a WCF protocol $\pi=(\pi_{1},\pi_{2})$ (where $\pi_{1}$ and $\pi_{2}$ are used instead of $\pi_{A}$ and $\pi_{B}$ to avoid confusion in subsequent protocol description) may find its stand-alone security insufficient for the upper-level task.
To illustrate, consider an attacking strategy for an unbalanced WCF protocol $\mu=(\mu_{A},\mu_{B})$ constructed from two instances of balanced WCF protocol $\pi=(\pi_{1},\pi_{2})$. The protocol $\mu$ aims to model the scenario where Alice and Bob are concurrently running two instances of WCF with an undetermined execution order. Assume the protocol $\pi$ follows the quantum comb model [39, 38] with $n+1$ (assume that $n$ is odd without loss of generality) rounds of communication, where Alice and Bob send messages alternately. Then $\pi_{1}$ and $\pi_{2}$ can be characterized by the ordered set $\mathcal{T}_{\pi}$ shown in Figure 5(b), where the subset $\{t_{i}\}_{i=0}^{n}$ is totally ordered. The comb model further requires that
@@
The details of these causal boxes are shown in Appendix A.3.
@@
We demonstrate that the stand-alone security of $\mu$ can be compromised even if that of $\pi$ is maintained. As per [29, Lemma 4], if $\pi$ is a $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol and $\epsilon<1/6$, $\mu$ is expected to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol. However, this expectation is not met as there exists a strategy for a dishonest Bob to win the game with probability $\frac{1}{2}$. Before elaborating on this strategy, we introduce the delay box, denoted by $\Theta_{f}$. This box essentially transforms a message at position $t$ to a future position $t^{\prime}$ according to some delay function $f$. For a rigorous definition of delay box, see Appendix A.3.
Consider the parallel composition of delay boxes $\theta=\theta_{1}\|\theta_{2}$ depicted in Figure 6. The delay functions are defined as
@@
Proposition 8 shows that Figure 6(a) emulates a WCF protocol between Alice and herself. See Appendix B.1 for the proof.
@@
where $x_{1}^{i}=x_{1}x_{2}\dots x_{i}$ and the result of the $i$th instance is denoted by $c_{A,i}$. Notably, inequality (26) is an implicit but crucial condition for the security of unbalanced WCF protocol introduced in [29].
@@ 4.2 Global security
What is the sufficient condition for Inequality (26) to hold? More broadly speaking, what is the safe way to compose WCF protocols? Despite negative results discussed in Section 3 on the possibility of abstracting WCF as a black-box abstraction, there are still certain ways of composition, satisfying Inequality (26). To explore this, we generalize Inequality (26), introduce the concept of global secrecy and specify the sufficient condition for it.
@@
The proof of Theorem 10 is in Appendix B.2. The proof idea is by contradiction. If the adversary can break the global security of the composed protocol, then the adversary can also break the stand-alone security of the WCF protocol, which contradicts the fact and thus completes the proof. Theorem 10 establishes the sufficient condition that the WCF protocols can be composed while maintaining the overall stand-alone security.
For more complicated functionalities, it is usually easy to construct asymmetric imperfect protocols in which one party can cheat more than the other party. Unbalanced WCF protocols make it possible for two parties to fairly choose one of these asymmetric imperfect protocols at random. This results in a less asymmetric imperfect protocol. Therefore, unbalanced WCF protocols play a key role in constructing optimal protocols of more complicated functionalities such as SCF and BC [29, 28]. An unbalanced $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol can be construct from balanced $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocols where $z=0.b_{1}b_{2}\dots b_{n}$ is in binary representation, as is shown in Figure 7.
@@
As an example, we apply Theorem 10 to this construction to prove the stand-alone security of the unbalanced WCF protocol in Corollary 11. The idea is to prove the stand-alone security of $\tau_{i}(\pi_{1}\|...\|\pi_{i})$ constructed from $\tau_{i-1}(\pi_{1}\|...\|\pi_{i-1})$ and $\pi_{i}$ using Theorem 10 recursively. The detailed proof can be found in Appendix B.3.
@@ A.1 Causal box
A cut is bounded if $\exists t\in\mathcal{T},\mathcal{C}\in\mathcal{T}^{\preceq t}$. The set of all cuts of $\mathcal{T}$ is denoted by $\mathfrak{C}(\mathcal{T})$. The set of all bounded cuts of $\mathcal{T}$ is denoted by ¯$\bar{\mathfrak{C}}(\mathcal{T})$.
@@
Using cuts and causality function, we can formally define causal box.
@@
$\forall\mathcal{C},\mathcal{D}\in\bar{\mathfrak{C}}(\mathcal{T}),\mathcal{C}\subseteq\mathcal{D}:\Phi^{\mathcal{C}}=\mathrm{tr}_{\mathcal{D}\setminus\mathcal{C}}\circ\Phi^{\mathcal{D}}$ (33)
@@
Alternatively, the causal box $\Phi$ can be described in Choi-Jamiolkowski (CJ) representation. To be precise, the Choi operator $R_{\Phi}^{\mathcal{C}}$ for the CPTP map $\Phi^{\mathcal{C}}$ is a sesquilinear semi-definite form on $\mathcal{F}_{Y}^{\mathcal{C}}\otimes\mathcal{F}_{X}^{\chi(\mathcal{C})}$ such that
$R_{\Phi}^{\mathcal{C}}(\psi_{Y}\otimes\psi_{X},\varphi_{Y}\otimes\varphi_{X})=\langle\psi_{Y}|\Phi^{\mathcal{C}}(|\bar{\psi}_{X}\rangle\langle\bar{\varphi}_{X}|)|\varphi_{Y}\rangle.$ (35)
@@ A.2 Connection of causal boxes
Definition 16 (Parallel composition of causal box).
@@
The above definition of parallel composition implicitly requires the causal boxes $\Phi$ and $\Psi$ share the same partially ordered set $\mathcal{T}$. However, a general causal box may not contain the information of the causal structure of some other causal box, that is, there is no such common set $\mathcal{T}$. In this case, to make the composition possible, we propose a more general version of parallel composition. Before that, we introduce the union of posets that might have intersection.
@@
Lemma 1920 and 21 suggest that Definition 18 is well defined.
@@
$\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}=\bigcup_{t\in\mathcal{C}}(\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}).$
@@
Assume $\chi_{1}$ and $\chi_{2}$ are causality functions for $\Phi$ and $\Psi$, respectively. Construct the function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ such that $\chi(\mathcal{C})=\chi_{1}(\mathcal{C}\cap\mathcal{T}_{1})\cup\chi_{2}(\mathcal{C}\cap\mathcal{T}_{2})$, then it is easy to verify the conditions in Definition 14. ∎
@@
For boxes $\Phi,\Psi,\Theta$ with compatible posets $\mathcal{T}_{\Phi},\mathcal{T}_{\Psi},\mathcal{T}_{\Theta}$,
@@
Let $\Phi=\{\Phi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{AB}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{CD}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be a causal box such that $\dim\mathcal{H}_{B}=\dim\mathcal{H}_{C}$ is its CJ representation. Let $\{|k_{C}\rangle\}$ and $\{|l_{C}\rangle\}$ be any orthonormal bases of $\mathcal{F}_{C}^{\mathcal{C}}$ and $\{|k_{B}\rangle\}$ and $\{|l_{B}\rangle\}$ be corresponding bases of $\mathcal{F}_{C}^{\mathcal{C}}$. The new box by looping from wire $B$ to wire $C$, denoted by $\Psi=\Phi^{C\hookrightarrow B}$, is a set of maps $\{\Psi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{A}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{D}^{\mathcal{C}})\}$ with CJ representation
@@ A.3 Delay box
$\theta_{1}$ is a delay box from input wire $C^{\prime}$ to output wire $C^{\prime\prime}$ and is derived from $f_{1}$ defined in Equation 23. Then, $U_{f_{1}}^{\mathcal{C}}$ is an isometry between $\mathcal{F}_{C^{\prime}}^{\chi_{f_{1}}(\mathcal{C})}$ and $\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}}$

### v3

@@ 3.2 Impossibility for a general resource
Composable security is defined by requiring that an actual protocol emulates an ideal resource within a certain accuracy. As established in Theorem 5, no actual protocol can approximate the ideal WCF resource defined in Definition 4. However, this impossibility result pertains only to this specific ideal WCF resource. By redefining the ideal WCF resource, potentially introducing additional structure, a protocol that securely constructs this modified ideal resource may still exist at the cost of a more complicated and less useful ideal WCF resource.
@@
  • Simple. Resource simplicity refers to abstraction of non-essential details of a protocol, thereby reducing its interaction complexity. This simplification facilitates the analysis of the resource in a more general context. Here we take the interaction complexity $\mathsf{Intcom}(S_{A})$, $\mathsf{Intcom}(S_{B})$ (defined in Section 2.2) to quantify it. Although the idea of a simple resource is rather subjective, we state a resource to be simple if its net interaction complexity ($\mathsf{Intcom}(S_{A})+\mathsf{Intcom}(S_{B})$) is of $\mathcal{O}(1)$.
Next, we will show that a fully expressing and simple WCF resource cannot be constructed from noiseless communication channels. Specifically, Theorem 6 states that, if a fully expressing resource for $\mathrm{WCF}(z,\epsilon)$ with small $\epsilon$ can be constructed from noiseless communication channels, then its interaction complexity cannot be small.
@@
To prove Theorem 6, we need a proposition rewritten from [32, Theorem 8.2].
@@
Some parts of the argument in the above proof are of more general interest. That is, if there exists a resource $\mathcal{S}$ that abstracts a two-party protocol $\pi$, it implies the existence of another protocol $\pi^{\prime}$ with the same interaction complexity as the resource $\mathcal{S}$. Furthermore, if the resource reduces the interaction complexity of $\pi$, then $\pi^{\prime}$ is a better protocol than $\pi$ in terms of interaction complexity. An example is delegated quantum computation (DQC), where the client aims to delegate computation to the server while maintaining privacy. The $S_{A}$ (against dishonest client) of DQC resource [43] abstracting the protocol in [44, 45] has no interaction on either side. This implies the existence of a simple DQC protocol with no interaction, wherein the client performs all the computation herself, ensuring both correctness and blindness.
@@ 4 Security of weak coin flipping under composition
Section 3 highlights the difficulties when formulating the security of WCF in the AC framework. Fortunately, despite these difficulties, there is a prospect of attaining positive outcomes regarding the composability of such protocols. This section starts with the limitation of the stand-alone security definition (Definition 3) in the context of composition. Specifically, for the construction of unbalanced WCF from a balanced one, we propose a condition (Inequality (34)) which, while implicit in [29], is essential in the security proof of the unbalanced WCF protocol depicted there. Subsequently, we introduce global security as an extension of the stand-alone definition to encompass possible interdependency between the outputs of the WCF protocol and other components in a larger system. Furthermore, we establish the condition under which global security is maintained.
@@ 4.1 Stand-alone security is not sufficient
For a cryptographic task, some security definitions seem fine from a stand-alone viewpoint, but may present some loopholes when composed with other protocols. For example, the accessible information criteria for quantum key distribution falls short of providing composable security due to the data-locking effect [2, 1]. A similar problem occurs in WCF. In particular, a WCF protocol $\pi=(\pi_{1},\pi_{2})$ (where $\pi_{1}$ and $\pi_{2}$ are used instead of $\pi_{A}$ and $\pi_{B}$ to avoid confusion in subsequent protocol description) may find its stand-alone security insufficient for upper-level tasks.
To illustrate, consider an attacking strategy for an unbalanced WCF protocol $\mu=(\mu_{A},\mu_{B})$ constructed from two instances of balanced WCF protocol $\pi=(\pi_{1},\pi_{2})$. The protocol $\mu$ aims to model the scenario where Alice and Bob are concurrently running two instances of WCF with an undetermined execution order. Assume the protocol $\pi$ follows the quantum comb model [41, 40] with $n+1$ (assume that $n$ is odd without loss of generality) rounds of communication, where Alice and Bob send messages alternately. Then, $\pi_{1}$ and $\pi_{2}$ can be characterized by the ordered set $\mathcal{T}_{\pi}$ shown in Figure 5(b), where the subset $\{t_{i}\}_{i=0}^{n}$ is totally ordered. The comb model further requires that
@@
The above conditions allow us to consider the wires in the two directions separately–one within the even subset of $\mathcal{T}_{\pi}$ and the other within the odd subset of $\mathcal{T}_{\pi}$.
@@
We demonstrate that the stand-alone security of $\mu$ can be compromised even if that of $\pi$ is maintained. As per [29, Lemma 4], if $\pi$ is a $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocol and $\epsilon<1/6$, $\mu$ is expected to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol. However, it fails to be a $\mathrm{WCF}(\frac{3}{4},\frac{3}{2}\epsilon)$ protocol, as there exists a strategy for dishonest Bob to win the game with probability $\frac{1}{2}$. Before elaborating on this strategy, we introduce the delay box, denoted by $\Theta_{f}$. This box essentially transforms a message at position $t$ to a future position $t^{\prime}$ according to some delay function $f$. For a rigorous definition of delay box, see Section A.3.
For simplicity, denote $\theta_{f_{1}},\theta_{f_{2}}$ as $\theta_{1},\theta_{2}$. Consider the parallel composition of delay boxes $\theta=\theta_{1}\|\theta_{2}$ depicted in Figure 6. The delay functions are defined as
@@
Proposition 8 shows that Figure 6(a) emulates a WCF protocol between Alice and herself. See Section B.1 for the proof.
@@
where $x_{1}^{i}=x_{1}x_{2}\dots x_{i}$ and the result of the $i$-th instance is denoted by $c_{A,i}$. Notably, inequality (34) is an implicit but crucial condition for the security of unbalanced WCF protocol introduced in [29].
@@ 4.2 Global security
What is the sufficient condition for Inequality (34) to hold? More broadly speaking, what is the safe way to compose WCF protocols? Despite negative results discussed in Section 3 on the possibility of abstracting WCF as a resource, there are still certain secure ways of composition, e.g., those satisfying Inequality (34). To explore this, we generalize Inequality (34), introduce the concept of global secrecy and specify the sufficient condition for it.
@@
The proof of Theorem 10 is in Section B.2. The proof idea is by contradiction. If the adversary can break the global security of the composed protocol, then the adversary can also break the stand-alone security of the WCF protocol, which contradicts the fact and thus completes the proof. Theorem 10 establishes the sufficient condition that the WCF protocols can be composed while maintaining the overall stand-alone security.
For more complicated functionalities, it is usually easy to construct asymmetric imperfect protocols in which one party can cheat more than the other party. Unbalanced WCF protocols make it possible for two parties to fairly choose one of these asymmetric imperfect protocols at random. This results in a less asymmetric imperfect protocol. Therefore, unbalanced WCF protocols play a key role in constructing optimal protocols of more complicated functionalities such as SCF and BC [29, 28]. An unbalanced $\mathrm{WCF}(z,\epsilon^{\prime})$ protocol can be constructed from balanced $\mathrm{WCF}(\frac{1}{2},\epsilon)$ protocols where $z=0.b_{1}b_{2}\dots b_{n}$ is in binary representation, as is shown in Figure 7.
@@
As an example, we apply Theorem 10 to this construction to prove the stand-alone security of the unbalanced WCF protocol in Corollary 11. The idea is to prove the stand-alone security of $\tau_{i}(\pi_{1}\|...\|\pi_{i})$ constructed from $\tau_{i-1}(\pi_{1}\|...\|\pi_{i-1})$ and $\pi_{i}$ using Theorem 10 recursively. The detailed proof can be found in Section B.3.
@@
5 Discussion
Our work demonstrates that the standalone security of protocols does not guarantee the security under certain types of composition. In particular, although quantum WCF protocol can be secure in the standalone scenario, it fails to be composably secure in the abstract cryptography framework. Moreover, when we construct new protocols with quantum WCF protocol, it remains secure in sequential composition, but exhibits security loopholes in concurrent composition. Similar security loopholes also arise in other two-party cryptographic tasks, such as bit commitment [25, 37]. For these tasks, though we have hoped for simple and good abstractions for actual protocols in the abstract framework, our work suggests that such abstraction does not exist, which makes security analysis of such protocols notably challenging.
@@ A.1 Causal box
A cut is bounded if $\exists t\in\mathcal{T},\mathcal{C}\subseteq\mathcal{T}^{\preceq t}$. The set of all cuts of $\mathcal{T}$ is denoted by $\mathfrak{C}(\mathcal{T})$. The set of all bounded cuts of $\mathcal{T}$ is denoted by ¯$\bar{\mathfrak{C}}(\mathcal{T})$.
@@
Using cuts and causality function, we can formally define the causal box.
@@
$\forall\mathcal{C},\mathcal{D}\in\bar{\mathfrak{C}}(\mathcal{T}),\mathcal{C}\subseteq\mathcal{D}:\Phi^{\mathcal{C}}=\mathrm{tr}_{\mathcal{D}\setminus\mathcal{C}}\circ\Phi^{\mathcal{D}},$ (41)
@@
Alternatively, the causal box $\Phi$ can be described in Choi-Jamiołkowski (CJ) representation. To be precise, the Choi operator $R_{\Phi}^{\mathcal{C}}$ for the CPTP map $\Phi^{\mathcal{C}}$ is a sesquilinear semi-definite form on $\mathcal{F}_{Y}^{\mathcal{C}}\otimes\mathcal{F}_{X}^{\chi(\mathcal{C})}$ such that
$R_{\Phi}^{\mathcal{C}}(\psi_{Y}\otimes\psi_{X},\varphi_{Y}\otimes\varphi_{X})=\langle\psi_{Y}|\Phi^{\mathcal{C}}\left(|\bar{\psi}_{X}\rangle\langle\bar{\varphi}_{X}|\right)|\varphi_{Y}\rangle.$ (43)
@@ A.2 Connection of causal boxes
Definition 16 (Parallel composition of causal boxes).
@@
The above definition of parallel composition implicitly requires the causal boxes $\Phi$ and $\Psi$ share the same partially ordered set $\mathcal{T}$. However, a general causal box may not contain the information of the causal structure of some other causal box. Then, a predefined common poset $\mathcal{T}$ may not exist. In this case, to make the composition possible, we propose a more general version of parallel composition. Before that, we introduce the union of posets that might have intersection.
@@
Lemma 1920 and 21 suggest that Definition 18 is well-defined.
@@
$\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{T}_{1}=\bigcup_{t\in\mathcal{C}}(\mathcal{T}^{\preceq t}\cap\mathcal{T}_{1}).$ (48)
@@
Assume $\chi_{1}$ and $\chi_{2}$ are causality functions for $\Phi$ and $\Psi$, respectively. Construct the function $\chi:\mathfrak{C}(\mathcal{T})\to\mathfrak{C}(\mathcal{T})$ as $\chi(\mathcal{C})=\chi_{1}(\mathcal{C}\cap\mathcal{T}_{1})\cup\chi_{2}(\mathcal{C}\cap\mathcal{T}_{2})$, then it is easy to verify the conditions in Definition 14. ∎
@@
For boxes $\Phi,\Psi,\Theta$ with compatible posets $\mathcal{T}_{\Phi},\mathcal{T}_{\Psi},\mathcal{T}_{\Theta}$, we have
@@
Let $\Phi=\{\Phi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{AB}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{CD}^{\mathcal{C}})\}_{\mathcal{C}\in\bar{\mathfrak{C}}(\mathcal{T})}$ be a causal box such that $\dim\mathcal{H}_{B}=\dim\mathcal{H}_{C}$ is its CJ representation. Let $\{|k_{C}\rangle\}$ and $\{|l_{C}\rangle\}$ be any orthonormal bases of $\mathcal{F}_{C}^{\mathcal{C}}$ and $\{|k_{B}\rangle\}$ and $\{|l_{B}\rangle\}$ be corresponding bases of $\mathcal{F}_{B}^{\mathcal{C}}$. The new box by looping from wire $B$ to wire $C$, denoted by $\Psi=\Phi^{C\hookrightarrow B}$, is a set of maps $\{\Psi^{\mathcal{C}}:\mathfrak{T}(\mathcal{F}_{A}^{\mathcal{T}})\to\mathfrak{T}(\mathcal{F}_{D}^{\mathcal{C}})\}$ with CJ representation
@@
Note that, to define a loop, it is necessary to specify two bases for each of the connected wire. With different bases for the wires, the result of the loop can be different.
@@ A.3 Delay box
$\theta_{1}$ is a delay box from input wire $C^{\prime}$ to output wire $C^{\prime\prime}$ and is derived from $f_{1}$ defined in Equation 31. Then, $U_{f_{1}}^{\mathcal{C}}$ is an isometry between $\mathcal{F}_{C^{\prime}}^{\chi_{f_{1}}(\mathcal{C})}$ and $\mathcal{F}_{C^{\prime\prime}}^{\mathcal{C}\cap\mathcal{T}_{\pi^{\prime}}}$.